# Fundamental thresholds for computational and erasure errors via the coherent information

### v1

@@ IV.3.3 Coherent information for depolarizing noise and erasure errors
$I_{l}=(k-b_{l}-2c_{l})\log 2-\overline{\log\left[\dfrac{\sum_{D_{l}}\mathcal{Z}_{D_{l}}}{\mathcal{Z}_{0}}\right]},$ (68)
@@
$I=\left(k-\langle b_{l}\rangle-2\langle c_{l}\rangle\right)\log 2-\Biggr{\langle}\overline{\log\left[\dfrac{\sum_{D}\mathcal{Z}_{D,l}}{\mathcal{Z}_{0,l}}\right]}\Biggr{\rangle}.$ (69)
@@
We can clearly distinguish two sources of for the reduction of CI: (i) by means of the erasure errors the environment can effectively measure the logical operators and directly reduce the CI by integer multiples of $\log 2$. (ii) The computational errors act on the remaining physical qubits, with their effects quantified by the cost in free energy of inserting domain walls in the equilibrium state, as given by the partition function $\mathcal{Z}_{0}$ Dennis et al. (2002); Wang et al. (2003). The erasure errors drive the transition to a fully symmetric state by removing interactions locally, thereby lowering the cost of inserting domain walls. This picture is general and does not make any assumptions about the underlying geometry and spatial structure of the QEC code. Moreover, since the details of the code enter via the variables $P^{X}_{\ell}$ and $P^{Z}_{\ell}$ defined in Eq. (37), the classical spin models derived are the same with and without erasure errors, except for the missing links in the latter. For simplicity, we assume identical depolarizing channels on each qubit, but this could also be relaxed by allowing qubit-dependent couplings $J_{x},J_{z},J_{1}$ and considering spatially correlated noise Chubb and Flammia (2021).
@@ V.1 Erasure errors
We show in Fig. 8 the CI for the surface code and the 4.8.8 color code. The optimal threshold problem in the 2D toric code can be mapped to a bond percolation problem on a square lattice Stace et al. (2009), which yields a threshold of $e_{\text{th}}=0.5$, saturating the fundamental limit set by the no-cloning theorem. We find the same threshold in our calculation with four-digit precision, as shown in Fig. 8. In terms of percolation theory, the CI can be viewed as the probability of the appearance of percolating clusters. At low enough erasure probability $e$, both logical string operators still have support on non-lost qubits in the lattice, and effectively can traverse the lattice. In contrast, at high $e$, they become confined within small, disconnected regions and therefore stop being well-defined. Therefore, we expect the CI and the percolating cluster probability to have the same scaling exponent. In fact, we find $\nu=1.33(4)$ and $\nu=1.4(1)$ for the surface code and the color code, respectively, which is, within statistical error bars, consistent with the $\nu=4/3\approx 1.33$ predicted by percolation theory Stauffer and Aharony (1992). A similar result for the toric code in the context of mixed-state topological order was obtained in Ref. Kuno et al. (2024). Furthermore, since both codes have the same threshold and the same scaling exponent, we observe further evidence that the 2D toric and color codes have the same thresholds against quantum erasures in the code capacity setting Dennis et al. (2002); Bombin et al. (2012).
@@ V.2 Computational and erasure errors
It is important to note that in Ref. Stace et al. (2009) the authors show a mapping for the 2D toric code under bit-flip noise and erasure errors in which the probability of ferromagnetic and anti-ferromagnetic links is renormalized by the presence of the erasures. This statistical mechanics model yields the same optimal threshold as the diluted RBIM, according to Ref. Ohzeki (2012). However, to the best of our knowledge, the two models are not directly related. At this point, we argue that the CI allows us to unambiguously derive the statistical mechanics mappings with the certainty that they must be correct as long as they faithfully reproduce the CI of the noisy code density matrix. In this sense, we propose the CI as a means to validate the statistical mechanics mappings derived using the approach in Refs. Dennis et al. (2002); Ohzeki (2012); Stace and Barrett (2010), and we conclude that the diluted RBIM  faithfully captures the optimal decoding phase transition in the 2D toric code under bit-flip noise and erasure errors.
@@
VI Conclusions
@@
Third, we rigorously show how erasure errors modify the statistical mechanics mappings associated with computational errors. The key insight is that qubit erasures can be regarded as fully depolarizing channels acting at known positions. Hence, the mappings derived for computational errors are extended by adding an average over missing bonds, in addition to the bond disorder produced by the computational errors. This intuition also applies to mappings in the stabilizer configuration picture Fan et al. (2024); Hauser et al. (2024); Lyons (2024). There, a fully depolarizing channel translates to an infinite coupling, locking neighboring spins rigidly together (see Appendix C for details). Similarly, one can use the CI to establish connections between families of statistical mechanics models and to validate existing QEC mappings that lack closed-form expressions Dennis et al. (2002); Stace and Barrett (2010); Stace et al. (2009).
There are several routes to further explore erasure errors and the CI-based methodology to determine fundamental thresholds of QEC codes. In the absence of computational errors, one can now estimate optimal thresholds very reliably, making it feasible to extend this approach beyond 2D topological codes to higher dimensional codes or low-density parity-check (LDPC) codes. In both cases, erasure errors have not been extensively studied, and the equivalence of a QEC code graph and associated percolation picture is not always exact Delfosse and Zémor (2013); Kovalev and Pryadko (2013); Woolls and Pryadko (2022).
In addition, one could study how the mappings are modified by coherent and/or correlated errors. The CI methodology for deriving statistical mechanics models can also be straightforwardly extended to LDPC codes, where statistical mechanics mappings have yet to be more broadly explored. Beyond the QEC perspective, mappings derived from quantum LDPC codes may also open the door to new classes of ordered states and topological phenomena beyond the paradigm of Euclidean space Placke et al. (2024); Roeck et al. (2024).
@@ Acknowledgements.
We thank Davide Vodola for useful discussions. L.C. and M.M. gratefully acknowledge funding by the U.S. ARO Grant No. W911NF-21-1-0007. M.M. furthermore acknowledges funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101114305 (“MILLENION-SGA1” EU Project), and this research is also part of the Munich Quantum Valley (K-8), which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. M.M. acknowledges funding from the ERC Starting Grant QNets through Grant No. 804247, and from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). L.C. and M.M. also acknowledge support for the research that was sponsored by IARPA and the Army Research Office, under the Entangled Logical Qubits program, and was accomplished under Cooperative Agreement Number W911NF-23-2-0216. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of IARPA, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. M.M. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 390534769. S.K. is supported by the National Research Foundation of Korea under grant NRF-2021R1A2C1092701 funded by the Korean government (MEST) and by the Institute of Information & Communication Technology Planning & Evaluation grant funded by the Korean government (Ministry of Science and ICT) (IITP-2024-RS-2024-00437191). SK would like to thank IQI (RWTH Aachen University) and the Forschungszentrum Jülich for the support for the visit to the IQI in the initial phase of this work. We used the python libraries LDPC Roffe (2022), NumPy Harris et al. (2020) and SciPy Virtanen et al. (2020).
@@ Appendix A Method for computing coherent information under erasure errors
Given a CSS stabilizer code $[[n,k,d]]$, there are $k$ pairs of logical operators $\{O^{z}_{i},O^{x}_{i}\}$ which are the generators of the logical group. Each operator can be interpreted as a binary vector of size $n$, where 0 (1) on position $j$ means empty (occupied) by either $X$ or $Z$ Pauli. Since the code is CSS we treat $X$ and $Z$ operators independently, although the same formalism can be applied to non-CSS codes. We also have parity check matrices $H^{x}$ and $H^{z}$ whose rows are the stabilizer generators, hence they have dimension $n\times(n-k)$. We define a loss vector $r_{l}$ of dimension $n$ which is all zeros except on the positions the erased qubit. The configuation of qubit erasures is denoted by the index $l=1,2,...,2^{n}$. In Refs. Amaro et al. (2020); Harris et al. (2018), the authors show an algorithm for determining whether the erasure $r_{l}$ has support on at least one representative of the logical operators $\{O^{z}_{i},O^{x}_{i}\}$ based only on $H^{z}$, $H^{x}$ and $r_{l}$. In other words, this indicates whether the logical operator was measured by the environment or not. If for the $i-$th logical qubit one of the two logical operators is measured then it becomes a logical bit. If both logical operators are measured then we call it a lost qubit. Computing the coherent information for a fixed erasure configuration $l$ means computing how many logical qubits, logical bits and lost logical qubits there are. Another way to interpret the appearance of logical bits and lost logical qubits is as a reduction in the number of generators of the logical group. If one or more logical operators cannot be recovered, then the logical group will have fewer generators than the original $2k$ logical generators. This affects the size of the code space, thereby decreasing the amount of information that can be protected by the QEC code.

### v2

@@ IV.3.3 Coherent information for Pauli noise and erasure errors
$I_{l}=k-b_{l}-2c_{l}-\overline{\log\left[\dfrac{\sum_{D_{l}}\mathcal{Z}_{D_{l}}}{\mathcal{Z}_{0}}\right]},$ (69)
@@
$I=k-\langle b_{l}\rangle-2\langle c_{l}\rangle-\Biggr{\langle}\overline{\log\left[\dfrac{\sum_{D}\mathcal{Z}_{D,l}}{\mathcal{Z}_{0,l}}\right]}\Biggr{\rangle}.$ (70)
@@
We can clearly distinguish two sources for the reduction of CI: (i) by means of the erasure errors the environment can effectively measure the logical operators and directly reduce the CI by integer multiples of $\log 2=1$. (ii) The computational errors act on the remaining physical qubits, with their effects quantified by the cost in free energy of inserting domain walls in the equilibrium state, as given by the partition function $\mathcal{Z}_{0}$ Dennis et al. (2002); Wang et al. (2003). The erasure errors drive the transition to a fully symmetric state by removing interactions locally, thereby lowering the cost of inserting domain walls. This picture is general and does not make any assumptions about the underlying geometry and spatial structure of the QEC code. Moreover, since the details of the code enter via the variables $P^{X}_{\ell}$ and $P^{Z}_{\ell}$ defined in Eq. (38), the classical spin models derived are the same with and without erasure errors, except for the missing links in the latter. For simplicity, we assume identical Pauli channels on each qubit, but this could also be relaxed by allowing qubit-dependent couplings $J_{x},J_{z},J_{1}$ and considering spatially correlated noise Chubb and Flammia (2021).
@@ V.1 Erasure errors
We show in Fig. 8 the CI for the surface code and the 4.8.8 color code. The optimal threshold problem in the 2D toric code can be mapped to a bond percolation problem on a square lattice Stace et al. (2009), which yields a threshold of $e_{\text{th}}=0.5$, saturating the fundamental limit set by the no-cloning theorem. We find the same threshold in our calculation with four-digit precision, as shown in Fig. 8. In terms of percolation theory, the CI, after renormalized and rescaled properly, can be viewed as the probability of the appearance of percolating clusters. At low enough erasure probability $e$, both logical string operators still have support on non-lost qubits in the lattice, and effectively can traverse the lattice. In contrast, at high $e$, they become confined within small, disconnected regions and therefore stop being well-defined. Therefore, we expect the CI and the percolating cluster probability to have the same scaling exponent. In fact, we find $\nu=1.33(4)$ and $\nu=1.4(1)$ for the surface code and the color code, respectively, which is, within statistical error bars, consistent with the $\nu=4/3\approx 1.33$ predicted by percolation theory Stauffer and Aharony (1992). A similar result for the toric code in the context of mixed-state topological order was obtained in Ref. Kuno et al. (2024). Furthermore, since both codes have the same threshold and the same scaling exponent, we observe further evidence that the 2D toric and color codes have the same thresholds against quantum erasures in the code capacity setting Dennis et al. (2002); Bombin et al. (2012).
@@ V.2 Computational and erasure errors
It is important to note that in Ref. Stace et al. (2009) the authors show a mapping for the 2D toric code under bit-flip noise and erasure errors in which the probability of ferromagnetic and anti-ferromagnetic links is renormalized by the presence of the erasures. This statistical mechanics model yields the same optimal threshold as the diluted RBIM, according to Ref. Ohzeki (2012). However, to the best of our knowledge, the two models are not directly related. At this point, we argue that the CI allows us to unambiguously derive the statistical mechanics mappings with the certainty that they must be correct as long as they faithfully reproduce the CI of the noisy code density matrix. In this sense, we propose the CI as a means to validate the statistical mechanics mappings derived using the approach in Refs. Dennis et al. (2002); Ohzeki (2012); Stace and Barrett (2010), and we conclude that the diluted RBIM  faithfully captures the optimal decoding phase transition in the 2D toric code under bit-flip noise and erasure errors for arbitrary code sizes. It is worth to clarify that the latter does not exclude that some statistical mechanics models, which are not directly related to the one derived from the CI, can represent the optimal decoding phase transition only in the thermodynamic limit. In other words, the advantage in deriving the statistical mechanics models from the CI is the built-in guarantee in representing the exact optimal decoding problem for arbitrary code sizes.
@@
VI Lift-connected surface code
Figure 11: Lift-connected surface (LCS) code with parameters $L=4$ and $\ell=1$. The stabilizers associated to the $\ell+1$ surface code sheet $q=0$ are depicted. There are four stabilizers per surface code sheet. Each $X$ ($Z$) stabilizer is constructed by taking the qubits of the respective $\ell+1$ surface code plus the left (right) qubit on the $q-1$ sheet and the lower (upper) qubit on the $q+1$ sheet. The support of $X_{L}$ and $Z_{L}$ logical operators associated to the $q=0$ surface code sheet is shown highlighted in purple. The other stabilizers and logical operators can be visualized by rotating the indices $q\rightarrow(q+1)\mod L$.
Figure 12: Depiction of $\sigma$ and $\tau$ spins and their location with respect to the qubits for the [[15,3,3]] lift-connected surface code. There are two $\sigma$ and $\tau$ spins on each surface code sheet. The weight-three spin interactions for the $\sigma$ spins are shown as green continuous, dashed and dotted lines. The third spin always comes from the $q+1$ surface code sheet on the right. See Appendix C for more details on the spin model.
In order to test our method beyond the realm of topological codes, we now study the lift-connected surface (LCS) code introduced in Ref. Old et al. (2024). To construct the code family, one takes two repetition codes of size $\ell+1$ and applies the lift-product Panteleev and Kalachev (2021, 2022) with circulant matrices of size $L$. As explained in Ref. Old et al. (2024), this procedure yields an LDPC code $[[n,k,d]]$ with parameters
$n=[(\ell+1)^{2}+\ell^{2}]L$
$k=L$
$d=\min(L,2\ell+1).$ (73)
This code is more easily interpreted as $L$ unrotated surface code sheets of linear size $\ell+1$ arranged in a periodic stack; see Fig. 11. Each stabilizer of the standard surface code is now extended over neighboring surface code sheets, leading to weight-six $X$ and $Z$ stabilizers in the bulk and weight-four stabilizers at the boundary. The connectivity needed for measuring stabilizers is thus 3D-local. The logical operators are extended over multiple surface code sheets, and their minimum-weight representatives lie along the diagonals of each surface code sheet. Let us note that, since the logical operators are not defined on the boundaries, these are not topological codes in the same sense as, for example, 3D surface codes.
The two parameters $(L,\ell)$ are independent; therefore, it is convenient to choose a line in the $(L,\ell)$ plane to study thermodynamic properties. Here, we choose $L=2\ell+1$, which yields an ever-growing code distance $d$ and the minimum number of physical qubits for a fixed distance Old et al. (2024). Other choices that keep either $L$ or $\ell$ constant result in a bounded code distance $d$ (see Eq. (VI)) which prevents the development of a finite threshold. The downside of the family $L=2\ell+1$ is its vanishing rate, $R=k/n=[(\ell+1)^{2}+\ell^{2}]^{-1}$, as $\ell$ grows. We leave the study of finite-rate LDPC codes for future works.
In the remainder of this section, we present numerical calculations of the coherent information (CI) for the LCS code under two noise models: erasure errors only, and a combination of computational and erasure errors. For the former, we provide results up to $d=17$, whereas for the latter, we are limited to $d=3$ and $d=1$. We also present, for the first time, the exact statistical mechanics mappings for LCS codes under bit/phase-flip and depolarizing noise, along with their direct extensions to include erasure errors. These mappings are used to compute the CI exactly in the case of combined bit/phase-flip and erasure errors. For depolarizing noise, we sample over erasure errors while treating the computational errors exactly.
VI.1 Erasure errors
Figure 13: Coherent information of lift-connected surface code under erasure errors. We obtain a threshold of $e_{th}=0.5$ and a critical exponent $\nu\approx 0.75$, which belongs neither to the 3D nor the 2D percolation universality class Stauffer and Aharony (1992).
In Fig. 13, we show the CI for the LCS code under erasure errors. We find that the optimal threshold is located at $e_{\text{th}}=0.5$, which saturates the bound imposed by the no-cloning theorem Bennett et al. (1997). The exponent $\nu\approx 0.75$ does not match the exponents of 2D or 3D percolation Stauffer and Aharony (1992), reflecting the more complex nature of the stabilizers and logical operators compared to those in 2D and 3D topological codes.
VI.2 Computational and erasure errors
Figure 14: Pseudo-threshold, as crossing of the CI between $d=3$ and $d=1$ (single qubit) for various erasure probabilities $e$ and computational error probabilities $p$. In blue and red we show the pseudothreshold of the LCS surface code and the standard surface code, respectively. Both codes have very similar pseudo-thresholds under the same error processes.
Let us now introduce the statistical mechanics mappings for the LCS codes. First, we observe that in the bulk, the stabilizers take the form $P^{X}_{\ell}=\sigma_{i,q}\sigma_{j,q}\sigma_{i,q+1}$ and $P^{Z}_{\ell}=\tau_{n,q}\tau_{m,q}\tau_{m,q+1}$, where the indices $(i,q)$ denote the 2D coordinates on the $q$-th surface code sheet. In other words, the two-spin terms within each surface code sheet are now lifted to three-spin interactions. The newly added spins are always at the same position $i$, but on neighboring surface code sheets. These three-spin terms are depicted in Fig. 12 for the example of a $[[15,3,3]]$ LCS code. The Hamiltonian is then written as
$H_{LCS}=\sum_{\langle i,j\rangle,\langle n,m\rangle}\eta^{x}_{ij}\left(J_{x}-\dfrac{J_{1}}{2}\right)\dfrac{\sigma_{i,q}\sigma_{j,q}\sigma_{i,q+1}}{2}$
$+\eta^{z}_{nm}\left(J_{z}-\dfrac{J_{1}}{2}\right)\dfrac{\tau_{n,q}\tau_{m,q}\tau_{m,q+1}}{2}$
$+\eta^{z}_{nm}\eta^{x}_{ij}J_{1}\dfrac{\sigma_{i,q}\sigma_{j,q}\sigma_{i,q+1}\tau_{n,q}\tau_{m,q}\tau_{m,q+1}}{4}.$ (74)
Importantly, for the qubits located at the boundaries, some terms in the Hamiltonian will involve one- and two-spin interactions; see Appendix C for the exact description of the $[[15,3,3]]$ Hamiltonian. To the best of our knowledge, this family of statistical mechanics models has not been studied in the literature. Notably, Eq. (VI.2) holds for any choice of $(L,\ell)$ and is therefore not limited to our case study of $L=2\ell+1$.
In Fig. 14, we show a phase diagram for finite erasure and computational error probabilities (bit/phase-flip and depolarizing noise), defined by the crossing of the $d=3$ and $d=1$ CI curves. Remarkably, the pseudo-thresholds of the LCS code are very close to those of the rotated surface code studied in Sec. V. Together with the $50\%$ threshold found for erasures only, we note that the LCS code, in the code capacity setting, inherits most of the error correction capabilities of the 2D standard surface code, despite having a growing number of logical qubits.
VII Conclusions
@@
Third, we rigorously show how erasure errors modify the statistical mechanics mappings associated with computational errors. The key insight is that qubit erasures can be regarded as fully depolarizing channels acting at known positions. Hence, the mappings derived for computational errors are extended by adding an average over missing bonds, in addition to the bond disorder produced by the computational errors. This intuition also applies to mappings in the stabilizer configuration picture Fan et al. (2024); Hauser et al. (2024); Lyons (2024). There, a fully depolarizing channel translates to an infinite coupling, locking neighboring spins rigidly together (see Appendix D for details). Similarly, one can use the CI to establish connections between families of statistical mechanics models and to validate existing QEC mappings that lack closed-form expressions Dennis et al. (2002); Stace and Barrett (2010); Stace et al. (2009).
We also showcase the generality of the CI approach by studying optimal thresholds for a low-density parity-check (LDPC) code: the lift-connected surface code Old et al. (2024). This code has a growing number of logical qubits, making it the first code with this feature to be systematically studied using CI. Under erasure errors only, we find a $50\%$ optimal threshold and a critical exponent that, so far, has not been related to any known percolation transition. Furthermore, we derive a family of statistical mechanics mappings describing the optimal decoding problem for this code and identify a 3D-local model exhibiting chiral spin interactions. An in-depth study of the statistical physics models arising from this code, potentially involving larger system sizes, is left for future work. Finally, we compute the CI numerically and find pseudo-thresholds under combined computational and erasure errors that are comparable to those of the standard surface code. This indicates that the lift-connected surface code offers high levels of noise protection, despite having a growing number of logical qubits.
There are several routes to further explore erasure errors and the CI-based methodology to determine fundamental thresholds of QEC codes. In the absence of computational errors, one can now estimate optimal thresholds very reliably, making it feasible to extend this approach beyond 2D topological codes to higher dimensional codes or LDPC codes with finite-high rate. In both cases, erasure errors have not been extensively studied, and the equivalence of a QEC code graph and associated percolation picture is not always exact Delfosse and Zémor (2013); Kovalev and Pryadko (2013); Woolls and Pryadko (2022).
In addition, one could study how the mappings are modified by coherent and/or correlated errors. A point of great interest is the exploration of new families of statistical mechanics models that can be obtained from LDPC codes with different features (e.g., finite rate), potentially uncovering new physical phenomena arising from exotic QEC codes. Beyond the QEC perspective, mappings derived from quantum LDPC codes may also lead to new classes of ordered states and topological phenomena beyond the paradigm of Euclidean space Placke et al. (2024); Roeck et al. (2024).
@@ Acknowledgements.
We thank Davide Vodola for useful discussions. We thank Josias Old for fruitful discussions about the lift-connected surface code. L.C. and M.M. gratefully acknowledge funding by the U.S. ARO Grant No. W911NF-21-1-0007. M.M. furthermore acknowledges funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101114305 (“MILLENION-SGA1” EU Project), and this research is also part of the Munich Quantum Valley (K-8), which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. M.M. acknowledges funding from the ERC Starting Grant QNets through Grant No. 804247, and from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). L.C. and M.M. also acknowledge support for the research that was sponsored by IARPA and the Army Research Office, under the Entangled Logical Qubits program, and was accomplished under Cooperative Agreement Number W911NF-23-2-0216. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of IARPA, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. M.M. acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 390534769. S.K. is supported by the National Research Foundation of Korea under grant NRF-2021R1A2C1092701 funded by the Korean government (MEST) and by the Institute of Information & Communication Technology Planning & Evaluation grant funded by the Korean government (Ministry of Science and ICT) (IITP-2024-RS-2024-00437191). SK would like to thank IQI (RWTH Aachen University) and the Forschungszentrum Jülich for the support for the visit to the IQI in the initial phase of this work. The authors gratefully acknowledge the computing time provided to them at the NHR Center NHR4CES at RWTH Aachen University (Project No. p0020074). This is funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for national high performance computing at universities. We used the python libraries LDPC Roffe (2022), NumPy Harris et al. (2020) and SciPy Virtanen et al. (2020).
@@ Appendix A Method for computing coherent information under erasure errors
Given a CSS stabilizer code $[[n,k,d]]$, there are $k$ pairs of logical operators $\{O^{z}_{i},O^{x}_{i}\}$ which are the generators of the logical group. Each operator can be interpreted as a binary vector of size $n$, where 0 (1) on position $j$ means empty (occupied) by either $X$ or $Z$ Pauli. Since the code is CSS we treat $X$ and $Z$ operators independently, although the same formalism can be applied to non-CSS codes. We also have parity check matrices $H^{x}$ and $H^{z}$ whose rows are the stabilizer generators, hence they have dimension $n\times(n-k)$. We define a loss vector $r_{l}$ of dimension $n$ which is all zeros except on the positions of the erased qubit. The configuation of qubit erasures is denoted by the index $l=1,2,...,2^{n}$. In Refs. Amaro et al. (2020); Harris et al. (2018), the authors show an algorithm for determining whether the erasure $r_{l}$ has support on at least one representative of the logical operators $\{O^{z}_{i},O^{x}_{i}\}$ based only on $H^{z}$, $H^{x}$ and $r_{l}$. In other words, this indicates whether the logical operator was measured by the environment or not. If for the $i-$th logical qubit one of the two logical operators is measured then it becomes a logical bit. If both logical operators are measured then we call it a lost qubit. Computing the coherent information for a fixed erasure configuration $l$ means computing how many logical qubits, logical bits and lost logical qubits there are. Another way to interpret the appearance of logical bits and lost logical qubits is as a reduction in the number of generators of the logical group. If one or more logical operators cannot be recovered, then the logical group will have fewer generators than the original $2k$ logical generators. This affects the size of the code space, thereby decreasing the amount of information that can be protected by the QEC code.